# Automated Tabletop Exfoliation and Identification of Monolayer Graphene Flakes

### v1

@@ II.3 Wafer Imaging and Monolayer Identification
It is well known that the optical transmission of monolayer graphene is an approximately constant 97.7% for wavelengths in the visible range[26]. However, beginning with Geim and Novoselov’s work[10], most researchers exfoliate on silicon wafers with thin oxide to improve contrast, as the reflection from the underlying silicon roughly doubles apparent contrast, and thin-film interference effects from the oxide layer - which is typically about 90 or 300nm thick - further enhance visibility, enabling monolayer graphene to be seen through standard optical microscopes without heavy image processing.
@@
For the purpose of automating flake detection, it is beneficial to choose lighting conditions and camera settings to further enhance contrast. All imaging used the provided LED illumination of the Leica DM6 M and a 10x magnification objective (Leica Objective #566503, NA 0.3). In software, a flat-field correction was applied to all images in order to remove vignetting, and a suitable white-balance of the images was chosen and held constant through the study, resulting in all images being saved in false-color and enhancing monolayer contrast against the silicon background.
@@
Though all sample wafers have nominally uniform 300nm thick layers of silicon oxide, it was noted during the study that the manufacturer’s specified tolerance of 5% oxide thickness resulted in varying background colors within individual wafers and from wafer to wafer. Thus, it was necessary to determine empirical relations between apparent background and monolayer color across a range of backgrounds. These relations were initially determined by collecting over 200 monolayer-background color pairs from the output of a less accurate algorithm across several samples, and were then held fixed throughout the experiment.
@@
The computer vision (CV) algorithm detects the background color in each image, and uses the previously calculated empirical linear relations between background and monolayer colors in the R, G, and B channels under the established lighting settings to predict monolayer appearance. Clusters of pixels within a certain Euclidean distance of the predicted flake color were then identified and labeled, with the corresponding image saved for subsequent review.
@@
All images identified as containing monolayers were manually categorized to identify freestanding monolayers and eliminate false positives before being uploaded to a shared repository for use in device fabrication; a wafer-scale stitched image was generated for each sample with labeled flake locations to facilitate retrieval, as shown in Figure 2.
@@
As discussed in Results, the algorithm was less reliable in detecting monolayers with areas below 100 m - corresponding to a smallest length dimension of 5-10 pixels. However, the most desirable monolayers for assembling TBG stacks are typically above 400 m in area. We accordingly chose 200 m as the cutoff value above which we presumed all flakes were found. In addition to this dependence on area, other significant limitations of the algorithm include sensitivity to lighting settings and a failure to detect monolayers located on backgrounds outside the range used to determine the initial empirical relations. Prepared samples were excluded from the study if large discolored areas with no detected monolayer flakes were present while other regions of the sample exhibited a significant distribution of monolayer flakes, as this indicates a failure of the flake-finding algorithm due to background color. However, as the background color is not affected by any step of the exfoliation process (including anneals up to 105C), this is not expected to bias the reported results. For further details on the flake-finding algorithm, see Bittner et al. (2024), in preparation[25].
@@ III Results
In total, 33 sample wafers were created over a range of peel speeds and anneal temperatures, as shown in Table 2. All wafers were created by applying the populated tape with an average pressure kPa and peeling with an angle that repeatably swept from 60 to 80 across each wafer.
Based on the empirical observation that various manual exfoliation techniques seem to produce similar results, the parameter space for exfoliation appears to be fairly flat. That is, while different techniques and user experience may moderately increase flake yield per cm, it is expected that no dramatically optimal parameters will be found. Thus, it was deemed satisfactory to find a point in parameter space that produces a number of large, freestanding monolayers with low levels of polymer residue at a rate comparable to both manual exfoliation, and sufficient to feed the pipeline of device fabrication in the lab.
@@
After exfoliation, the CV flake searching algorithm was applied and the freestanding monolayers of area greater than 200 m ("useful" monolayers) tabulated, with average yields per wafer shown in Table 7 (see Appendix B).
As the graphite crystals used in this experiment were all formed by natural processes, it is expected that the size of crystal domains therein will vary, an expectation born out by cursory examinations of the distribution of flake areas, as exfoliated monolayer graphene flakes are typically single-crystal. Furthermore, repeatedly cleaving a flake via adhesion to tape during the population step naturally leads to a distribution of flake dimensions. This can be seen by considering a simple model of an -layer crystal as cleaving into fractions of thickness and for some positive constant . Using this model, upon repeated cleaving steps - as occur during tape population - the probability to find a flake of thickness after cleaves is given by , where values of represent monolayers with areas smaller than the starting crystal. This inevitable distribution of flake thickness (and monolayer areas) is further compounded by differences in flake thickness and crystal domain size in the starting graphite, and may account for some of the variance seen in the observed yields.
To account for this inherent variance in the manual population step, the useful monolayer counts were then normalized by the area of graphite remaining on the populated tape post-exfoliation as a proxy for the initial graphite area. As shown in Table 3, slower peel speeds tend to yield more useful monolayers.
| UML/A | 1 m/s | 10 m/s | 100 m/s | 1 cm/s | Mean |
| (N/mm) | |||||
| 25C | 1.70.1 | 2.41.6 | 1.40.9 | 0.71.2 | 1.61.3 |
| 65C | 3.40.8 | 2.9 | 2.61 | 1.8 | 2.80.9 |
| 105C | 2.9 | 0.9 | 1.6 | 2.2 | 1.60.9 |
@@
As shown in the preceding tables, the standard deviations of most measured metrics are comparable to the mean observed values; while this precludes strong conclusions about the dependence of yield on peel speed or anneal temperature, some weak trends may still be observed. This is clarified by calculating Pearson’s product-moment correlation coefficient , which describes the strength of linear associations between two variables - true positive (negative) linear dependence is represented by , with a total lack of dependence represented by . Here, while absolute values of indicate a weak relationship between peel speed and normalized UML yields, it is clear that an uncontrolled variable is strongly affecting actual results. We presume this to be the inherent variance of tapes manually populated with natural graphite, though other uncontrolled factors such as ambient humidity or substrate contamination may also contribute.
| UML/A | A/A | UML Yield | |||
| (N/wafer) | (N/mm) | (m/mm) | (%) | (m/flake) | |
@@
Though monolayer yield per wafer seems to somewhat increase at slower peel speeds and 65C, the increase in residue and added (unattended) time per exfoliation favors using higher peel speeds and lower temperature anneals. Aside from the annealing and peeling steps, typical sample preparation for the eXfoliator took approximately 15 minutes, with another 60 minutes post-exfoliation to scan and search each wafer. As the eXfoliator can be prepared and run in parallel with the mapping microscope, maximum sample production is in practice reached by all recipes that take less than 45 minutes to complete. With the approximate time to complete each recipe given by Table 12, this limits efficient exfoliation recipes to peel speeds of 100 m/s or higher. As there is no strong correlation between individual monolayer area and anneal temperature or peel speed, the figure of merit for this study is taken to be number of monolayers per starting area of graphite; as shown in Table 3, annealing at 65C seems to yield better results than the other temperatures. However, because tape adhesive is a viscous polymer that readily flows at these temperatures, slower peel speeds and higher anneal temperatures result in increased levels of polymer residue on the wafer, as shown in Figure 4. Monolayers are primarily cleaved from the bottom of larger flakes and are thus often pristine regardless of the amount of residue elsewhere on the wafer, but stamp-based stacking techniques are often hindered by the presence of thick tape residue nearby. We have no solvent-free technique for reliably removing heavy polymer residue left by Scotch magic tape without damaging flakes, so it is desirable to use exfoliation recipes that minimize tape residue while retaining sufficient yield of flakes. With the additional constraints of time and residue, the preferred exfoliation recipe explored in this study was determined to be annealing at 25C and peeling at 100 m/s.
Overall, approximately 35% of all monolayers detected with areas greater than 200 m were freestanding, useful monolayers. As shown in Appendix C, the total and useful monolayer yield per wafer of most parameter pairs did not significantly deviate from the average, suggesting that peel speed and anneal temperature are not the dominant causes of flake area distribution. The effect of uncontrolled variables, such as humidity and initial graphite thickness, was not examined in this study, but the eXfoliator design allows for further investigation if desired.
@@ IV Conclusions
This paper introduced a simple, compact device for the semi-automated exfoliation of monolayer materials, further describing the imaging system used to characterize the performance of the device across a range of parameters. Overall, the eXfoliator system successfully produces sufficient numbers of large, freestanding monolayers for use in device fabrication, with an output of approximately three large monolayers per wafer. The eXfoliator is believed to improve reproducibility and throughput compared to manual exfoliation, but further studies of flake yield with significantly more samples or a more reproducible technique for populating the initial tape are required. The eXfoliator retains accessibility, with new users typically requiring three or fewer sample preparations to reach near-maximum efficiency. As manual exfoliation may take months or years for a user to reach a similar rate of flake production, the eXfoliator should have significant utility in labs whose research intensively relies on high-quality graphene monolayers. Unlike Brookhaven’s QPress, the eXfoliator is not presently operated in an inert gas atmosphere, but its small footprint allows for placement into a glove-box if exfoliation of air-sensitive materials is desired.
The investigation of flake production as a function of anneal temperature and peel speed revealed that Scotch tape exfoliation does not strongly depend on these two factors at the parameter values investigated. In practice, the production rate of useful monolayers was most heavily affected by the total time to exfoliate each sample wafer and the amount of polymer tape residue remaining on an exfoliated wafer.
The eXfoliator’s design is well-suited for production of other van der Waals’ materials, such as hexagonal boron nitride and transition metal dichalcogenides, and allows for further exfoliation studies using other substrates, materials, and parameters. Current limitations of the eXfoliator design include its low cooling rate, placing a lower limit on the tape-silicon dwell time for elevated temperatures, and lack of an implemented rotational degree of freedom in the Y-Z plane, which prevents ’roll-on’ techniques to apply populated tape to silicon from being implemented.
@@ V Acknowledgments
The authors thank Chaitrali Duse and Steven Tran for their aid in preparing several samples; Joe Finney, Rupini Kamat, and Greg Zaborski Jr. for their insight and feedback in designing a machine to replicate manual exfoliation of monolayer graphene; Andy Mannix for useful discussions about the exfoliation process; and Feng Wang, Jingxu Xie, and Emma Regan for useful discussions about the reasoning behind their exfoliator’s design. Experimental measurements and analysis, parts for building the automated exfoliator, and acquisition of the motorized microscope for efficient imaging of the output of the exfoliator, were supported by the U.S. Department of Energy, Office of Basic Energy Sciences (DE-SC0021984). Supporting infrastructure was funded in part by the Gordon and Betty Moore Foundation through Grant No. GBMF3429.
@@ VI Appendix A - eXfoliator Parts
LCF300 25lb universal load cell, via FUTEK (FSH04274) with USB cable and output (FSH04720, FSH04742)
@@
We encountered a software bug in the firmware of the hotplate which made it insensitive to minor incremental changes in weight. Given that our design relied on making gradual contact with live feedback from the weight sensor, the hot plate could not be used for its originally intended purpose of providing both computerized temperature control and force feedback, inspiring the incorporation of a load cell.
@@ VII Appendix B - Additional Tables
Table 10 shows the percentage of all detected monolayers (including monolayers smaller than 200 m and those attached to bulk) that were useful monolayers, with slower peel speeds and an anneal temperature of 65C tending to yield a higher proportion of useful monolayers.

### v2

@@ II.3 Wafer Imaging and Monolayer Identification
It is well known that the optical transmission of monolayer graphene is an approximately constant 97.7% for wavelengths in the visible range[30]. However, beginning with Geim and Novoselov’s work[10], most researchers exfoliate on silicon wafers with thin oxide to improve contrast, as the reflection from the underlying silicon roughly doubles apparent contrast. In addition, thin-film interference effects from the oxide layer, typically chosen to be about 90 or 300nm thick, further enhance visibility, enabling monolayer graphene to be recognized in standard optical microscope images without heavy image processing.
@@
Choosing lighting conditions and camera settings to further enhance contrast can help in automating flake detection. All our imaging uses the provided LED illumination of the Leica DM6 M and a 10x magnification objective (Leica Objective #566503, NA 0.3). In software, we apply a flat-field correction to all images to remove vignetting, and we choose a suitable white-balance which we hold constant through the study. Thus all images are saved in false-color, with enhanced contrast for monolayers against the background appearance of the oxidized silicon.
@@
Though all sample wafers have nominally uniform 300nm thick layers of silicon oxide, we noted that the manufacturer’s specified tolerance of 5% oxide thickness resulted in varying background colors within individual wafers and from wafer to wafer. Thus, we had to determine empirical relations between apparent background and monolayer color across a range of backgrounds. To determine these relations we collected over 200 monolayer-background color pairs from the output of a less accurate algorithm across several samples. We held the relations fixed throughout the study.
@@
The computer vision (CV) algorithm detects the background color in each image, and uses the previously calculated empirical linear relations between background and monolayer colors in the R, G, and B channels under the established lighting settings to predict monolayer appearance. Clusters of pixels within a certain Euclidean distance of the predicted flake color are then identified and labeled, with the corresponding image saved for subsequent review.
@@
All images identified as containing monolayers are manually categorized to identify freestanding monolayers and eliminate false positives before being uploaded to a shared repository for use in device fabrication; a wafer-scale stitched image is generated for each sample. Flake locations are labeled to facilitate retrieval, as shown in Figure 2.
@@
As discussed in Results, the algorithm was less reliable in detecting monolayers with areas below 100 m2 - corresponding to a smallest length dimension of 5-10 pixels. However, the most desirable monolayers for assembling TBG stacks are typically above 400 m2 in area. We accordingly chose 200 m2 as the cutoff value above which we presumed all flakes were found. In addition to this dependence on area, other significant limitations of the algorithm include sensitivity to lighting settings and a failure to detect monolayers located on backgrounds outside the range used to determine the initial empirical relations. Prepared samples were excluded from the study if large discolored areas with no detected monolayer flakes were present while other regions of the sample exhibited a significant distribution of monolayer flakes, as this indicates a failure of the flake-finding algorithm due to background color. However, as the oxide thickness (and resulting background color) is not significantly affected by any step of the exfoliation process (including brief atmospheric anneals up to 105∘C), this is not expected to bias the reported results. For a similar, more sophisticated flake-finding algorithm, see Uslu et al., 2024[28].
@@ III Results
In total, we created 33 sample wafers over a range of peel speeds and anneal temperatures, as shown in Table 2. In each case we applied the populated tape with an average pressure kPa, and peeled with an angle that swept from 60∘ to 80∘ across each wafer.
This study was initially aimed to measure the dependence of monolayer yield on peel speed and application temperature, but it has become clear that yield’s dependence on uncontrolled variables such as the population of the initial tape, ambient humidity, and substrate contamination is comparable to its dependence on the controlled parameters. A full study of the influence of these parameters likely requires more precise control of at least the initial population step. Though the eXfoliator provides the minimum platform required for such an experiment, a full study is beyond the scope of this work.
In fact, a satisfying outcome is finding a point in the space of controlled parameters that produces large, freestanding monolayers at a rate comparable to manual exfoliation and sufficient to feed the pipeline of device fabrication in the lab, with low levels of polymer residue.
@@
After exfoliation, the CV flake searching algorithm was applied and the freestanding monolayers of area greater than 200 m2 (“useful” monolayers) tabulated, with average yields per wafer shown in Table 8 (see Appendix B).
As the graphite crystals used in this experiment were all formed by natural processes, the size of crystal domains therein might be expected to vary. However, even in the unlikely event that each layer in an initial graphite flake were a single crystal domain, repeatedly cleaving a flake via adhesion to tape during the population step would naturally lead to a distribution of flake dimensions. This can be seen by considering a simple model of an -layer crystal as cleaving into fractions of thickness and for some positive constant . Using this model, upon repeated cleaving steps - as occur during tape population - the probability to find a flake of thickness after cleaves is given by , where values of represent monolayers with areas smaller than the starting crystal. This inevitable distribution of flake thickness (and monolayer areas), further compounded by differences in flake thickness and crystal domain size in the starting graphite, may account for some of the variance seen in the observed yields.
To account for this inherent variance in the manual population step, the useful monolayer counts were then normalized by the area of graphite remaining on the populated tape post-exfoliation as a proxy for the initial graphite area. As shown in Table 3, slower peel speeds tend to yield a greater number of useful monolayers.
| UML/Agraphite | 1 m/s | 10 m/s | 100 m/s | 1 cm/s | Mean |
| (N/mm2) | |||||
| 25∘C | 1.70.1 | 2.41.6 | 1.40.9 | 0.71.2 | 1.61.3 |
| 65∘C | 3.40.8 | 2.9 | 2.61 | 1.8 | 2.80.9 |
| 105∘C | 2.9 | 0.9 | 1.6 | 2.2 | 1.60.9 |
@@
As shown in the preceding tables, the standard deviations of most measured metrics are comparable to the mean observed values; though this precludes strong conclusions about the dependence of yield on peel speed or anneal temperature, some weak trends may still be observed. This is clarified by calculating Pearson’s product-moment correlation coefficient , which describes the strength of linear associations between two variables - true positive (negative) linear dependence is represented by , with a total lack of dependence represented by . The absolute values of indicate a weak relationship between peel speed and normalized UML yield, with only 6% of the variation in yield explained by variation in peel speed. We suspect the remaining variation mainly derives from inherent variance in the properties of tapes manually populated with natural graphite, though other uncontrolled factors such as ambient humidity or substrate contamination may also contribute.
| UML/Agr | AUML/Agr | UML Yield | |||
| (N/wafer) | (N/mm2) | (m2/mm2) | (%) | (m2/flake) | |
@@
Though at slower peel speeds and 65∘C monolayer yield per wafer seems to somewhat increase, the concomitant increase in residue and added (unattended) time per exfoliation are unfavorable. Thus, going forward we would prefer higher peel speeds and lower temperature anneals. Aside from the annealing and peeling steps, typical sample preparation for the eXfoliator takes approximately 15 minutes, with another 60 minutes post-exfoliation to scan and search each wafer. For comparison, we surveyed our group members with varying levels of experience with manual exfoliation, asking them how long they took to exfoliate graphene over their preferred substrate area, and how long they took to then identify all useful monolayers. By naive scaling of the survey results, we conclude that a typical user would take between 25 and 150 minutes to manually exfoliate the area produced by a single run of the eXfoliator, followed by a manual searching time of 30 to 120 minutes.
In practice, as the eXfoliator can be prepared and run in parallel with the mapping microscope, maximum sample production is in practice reached by all recipes that take less than the 45 minutes the microscope needs to map the relevant area. With the approximate time to complete each recipe given by Table 12, this limits efficient exfoliation recipes to peel speeds of 100 m/s or higher. As there is no strong correlation between individual monolayer area and anneal temperature or peel speed, the figure of merit for this study is taken to be number of monolayers per starting area of graphite; as shown in Table 3, annealing at 65∘C seems to yield better results than the other temperatures. However, because tape adhesive is a viscous polymer that readily flows at these temperatures, slower peel speeds and higher anneal temperatures result in increased levels of polymer residue on the wafer, as shown in Figure 4. Monolayers are primarily cleaved from the bottom of larger flakes and are thus often pristine regardless of the amount of residue elsewhere on the wafer, but stamp-based stacking techniques are often hindered by the presence of thick tape residue nearby. We have no solvent-free technique for reliably removing heavy polymer residue left by Scotch magic tape without damaging flakes, so it is desirable to use exfoliation recipes that minimize tape residue while retaining sufficient yield of flakes. With the constraints of time and residue, among the recipes explored in this study we identified annealing at 25∘C and peeling at 100 m/s to be the most promising for future use.
Overall, approximately 34% of all monolayers detected with areas greater than 200 m2 were freestanding, useful monolayers. As shown in Appendix C, the total and useful monolayer yield per wafer of most parameter pairs did not significantly deviate from the average, suggesting peel speed and anneal temperature are not the dominant causes of flake area distribution. The effect of uncontrolled variables, such as humidity and initial graphite thickness, was not examined in this study, but the eXfoliator design allows for further investigation if desired.
@@ IV Conclusions
This paper introduces a simple, compact device for semi-automated exfoliation of monolayer materials, and describes the imaging system used to characterize the performance of the device across a range of parameters. Overall, the eXfoliator system successfully produces sufficient numbers of large, freestanding monolayers for use in device fabrication, with an output of approximately three large monolayers per wafer. The eXfoliator appears to improve reproducibility and throughput compared to manual exfoliation, but to robustly validate this assessment further studies of flake yield with significantly more samples or a more reproducible technique for populating the initial tape would be needed. The advantage of the eXfoliator may be primarily in allowing beginning users to achieve a rate of flake production they would otherwise take months to reach. Unlike Brookhaven’s QPress, the eXfoliator isn’t presently operated in an inert gas atmosphere, but its small footprint allows for placement into a glove-box if exfoliation of air-sensitive materials is desired. Compared to simpler automated peeling setups, the eXfoliator allows for control of peel angle and application pressure[22]. As the application pressure may affect substrate-crystal adhesion, especially in the presence of atmospheric adsorbates such as water and hydrocarbons, and the peel angle is likely to affect the ratio of shear to normal forces in the crystal during peeling, these parameters are deemed relevant to control.
We find that flake yield from Scotch tape exfoliation does not strongly depend on anneal temperature and peel speed over the range of parameter values investigated. In practice, the production rate of useful monolayers was most heavily affected by the total time to exfoliate each sample wafer and the amount of polymer tape residue remaining on an exfoliated wafer.
The eXfoliator’s design is well-suited for production of other van der Waals materials, such as hexagonal boron nitride and transition metal dichalcogenides, and allows for further exfoliation studies using other substrates, materials, and parameters. Limitations of the existing eXfoliator design which could be addressed include its slow cooling rate, placing a lower limit on the tape-silicon dwell time if heating is used, and lack of rotational control in the Y-Z plane, so populated tape cannot be ”rolled on” to silicon substrates.
@@ V Acknowledgments
The authors thank Chaitrali Duse and Steven Tran for their aid in preparing several samples; Joe Finney, Rupini Kamat, and Greg Zaborski Jr. for their insight and feedback in designing a machine to replicate manual exfoliation of monolayer graphene; Andy Mannix for useful discussions about the exfoliation process; and Feng Wang, Jingxu Xie, and Emma Regan for useful discussions about the reasoning behind their exfoliator’s design. Experimental measurements and analysis, parts for building the automated exfoliator, and acquisition of the motorized microscope for efficient imaging of the output of the exfoliator, were supported by the U.S. Department of Energy, Office of Basic Energy Sciences (DE-SC0021984). Supporting infrastructure was funded in part by the Gordon and Betty Moore Foundation through Grant No. GBMF3429 and Grant No. GBMF9460. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC., a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525.
The authors have no conflicts to disclose. The data that support the findings of this study are available from the corresponding author upon reasonable request. The authors contributed to this manuscript in the following ways - Elijah Courtney: Conceptualization (equal); Data Curation; Formal Analysis (lead); Investigation; Methodology (lead); Software (support); Original Draft Preparation; Review and Editing (equal). Mihir Pendharkar: Conceptualization (equal); Methodology (support); Review and Editing (equal). Nathan Bittner: Conceptualization (support); Software (lead); Review and Editing (equal). Aaron Sharpe: Conceptualization (support); Formal Analysis (support); Methodology (support); Review and Editing (equal). David Goldhaber-Gordon: Conceptualization (support); Methodology (support); Project Administration; Supervision; Review and Editing (equal).
@@ VI Appendix A - eXfoliator Parts
LCF300 25lb universal load cell, via FUTEK (FSH04274) with USB cable (FSH04720, FSH04742)
@@
The hot plate’s firmware was insensitive to changes in weight below 2g. The design relied on making gradual contact with live feedback from the weight sensor, so the hot plate could not be used for the intended purpose of providing both digital temperature control and force feedback, inspiring the load cell’s incorporation.
@@ VII Appendix B - Additional Tables
Table 10 shows the percentage of all detected monolayers (including monolayers smaller than 200 m2 and those attached to bulk) that were useful monolayers, with slower peel speeds and an anneal temperature of 65∘C tending to yield a higher proportion of useful monolayers.